QUESTION 4

  1. Spot rate is $1.68. Net profit or loss per unit is:

1.00000 points   

QUESTION 5

  1. This question and the following three questions use this information. One year ago, you sold a put option on 100,000 euros with an expiration date of one year. You received a premium on the put option of $.04 per unit. The exercise price was $1.22. Assume that one year ago, the spot rate of the euro was $1.20, the one-year forward rate exhibited a discount of 2%, and the one-year futures price was the same as the one-year forward rate. From one year ago to today, the euro depreciated against the dollar by 4 percent. Today the put option will be exercised (if it is feasible for the buyer to do so).

    The spot rate on the euro depreciated by 4%. What is the spot rate for the euro after it depreciates? Don't use units (e.g., don't use a dollar sign).

1.00000 points   

QUESTION 6

  1. What was the total loss on the put option? Again, don't use a dollar sign, and remember this is the total loss on all shares
Answer #1

Total premium received on put option = 100000 euros * $0.04 per unit = $4000

Spot rate of Euro Depreciated by 4%, therefore Depreciation on Euro = 1.2 * 4% = 0.048

Therefore Spot rate of Euro after 1 year i.e. after it depreciates = 1.2 - 0.048 = $1.152 i.e. 1.152

Calculation of total loss on put option

Exercise price = $1.22

Spot rate now = $1.152

The buyer of put will exercise the option since spot rate has reduced

Therefore total outflow on expiration of put option = (1.22-1.152) * 100000 = 0.068*100000 = 6800

Total loss = Outflow now - Premium received = 6800-4000 = $2800 i.e. 2800